# Eccentric features in the gravitational wave phase of dynamically formed black hole binaries

### v1

@@ Abstract
We study the gravitational wave (GW) phase shift arising from Rømer delay in binary black hole (BBH) mergers formed dynamically in three-body systems, where both the inner orbit of the merging binary and the outer orbit are eccentric. We provide a semi-analytical model and several analytical approximations that allow for fast evaluation of both the temporal evolution and the maximum value of the phase shift. The highest phase shifts occur when the binary merges close to the pericentre of the outer orbit, and can in this case be orders-of-magnitude larger compared to the circular limit. At high outer orbit eccentricities, the orbital curvature leaves distinct imprints onto the phase shift if the binary passes the outer pericentre during its inspiral. By comparing with phase-shifts measured in numerical chaotic 3-body scatterings, we show that our model accurately describes the observed phase of dynamically assembled binary systems in realistic astrophysical scenarios, providing a way to directly determine their formation channel via single GW observations.
@@ 1 Introduction
One very interesting measure of the formation environment is how the presence of a third body induces a phase shift onto the gravitational waveform in the observer frame. In such a scenario, the BBH evolves while its COM is being accelerated. The acceleration causes a time-dependent Doppler shift in the gravitational waveform, which in principle can be measured and used to probe the exact dynamical environment within which the binary lives as it merges. Therefore, looking for GW phase shifts, or more generally environmental effects, could be the key to help us tell apart the large suit of sub-channels, provided they are present in the observable GW bands.
@@
This generalisation is vital as it accurately represents the case for dynamically assembled BBH mergers (see e.g., Portegies Zwart & McMillan, 2000; Samsing et al., 2014b; Rodriguez et al., 2016a, b; Askar et al., 2017; Samsing & D’Orazio, 2018; Liu & Lai, 2018; Hoang et al., 2018; Trani et al., 2022). An example is depicted in Fig. 1. This shows the formation of a BBH inspiral and merger formed through a chaotic scattering process between three black holes. These types of interactions are known to frequently take place in GCs and other dense environments, and have especially been suggested to be among the most reliable ways for forming BBH mergers with residual eccentricity across the observable GW bands (Samsing, 2018). Moreover, including and considering outer orbit eccentricity is not expected to just add a small perturbation to the problem, as the maximum GW phase shift scales as $\propto R^{-2}$, where $R$ is the distance between the BBH and the third BH (Meiron et al., 2017; Samsing et al., 2024). This implies that if the BBH merges close to the pericentre of its outer orbit, the GW phase shift may increase dramatically compared to a circular outer orbit with the same period. The aim of this paper is to investigate the GW Doppler shift of an eccentric BBH on a eccentric orbit around a third BH, by quantifying effects that are unique to the outer orbit being eccentric.
@@
Figure 1: Example of a 3-body scattering that induces a highly eccentric BBH merger, where two are initially in a binary and the third is incoming. Scenarios like this particular one give rise to Doppler-shifted GW signals due to binary COM acceleration in the observer frame. In this example, we use $M_{1}=15M_{\odot}$, $M_{2}=5M_{\odot}$, and $M_{3}=15M_{\odot}$.
@@ 2 Theory
As described in Samsing et al. (2024),the GW Doppler shift arising from the acceleration on an outer orbit can equivalently be viewed as the GW phase shift from Rømer delay (RD) between the binary trajectory and some non-accelerating reference trajectory. In Fig. 2 we show our setup in this framework. We depict in turquoise the trajectory of an initially highly (inner-orbit) eccentric BBH (BH1 & BH2) that inspirals and merges on an (outer-orbit) eccentric orbit around BH3. We compare this to a fictional reference binary, shown in orange. This binary has exactly the same parameters (e.g. initial eccentricity, semi-major axis, time to coalescence, binary masses), but moves on a straight line with a constant velocity $v_{m}$ that is equal to the tangential velocity of the perturbed binary at merger. Tracing these two paths back in time from merger, their gradual deviation is what gives rise to the time-dependent Rømer delay. If the distance between the two trajectories at some time $t$ is given by $l(t)$ (red dashed line in Fig. 2), then their relative maximum Rømer delay is $\Delta t(t)=l(t)/c$. As shown in the bottom sketch of Fig. 2, the difference in arrival time between the GWs from the two trajectories induces a phase shift onto the gravitational waveform.
It is important to realise that the $\Delta t$ shown in Fig. 2 is the intrinsic time delay between the two trajectories. In reality, an observer only sees the projection of $\Delta t$ in their own line of sight (LOS). This translates into the observer dependent GW phase shift, $\delta{\phi}$, mathematically as,
$\delta{\phi}\approx d{\phi}_{\text{RD}}\times\cos(i)\sin(j).$ (1)
Here $\delta{\phi}_{\text{RD}}$ is the GW phase shift from Rømer Delay (RD) as inferred by a distant observer, $d{\phi}_{\text{RD}}$ is the intrinsic (or maximum possible) GW phase shift from Rømer Delay, $i$ is the angle in the plane of the ‘outer binary’ (the binary composed of (BH1,BH2)+BH3) between the LOS of the observer and the line connecting the location of BBH and the COM of the 3-body system, and $j$ is the angle between the LOS and the angular momentum vector of the ‘outer binary’. In this work, we solely focus on the intrinsic phase shift $d\phi_{\mathrm{RD}}$ and denote it simply by $d\phi$; factoring in the trivial observer dependent factor will be investigated in later work. Other sources of dephasing in gravitational waveforms in dynamically assembled systems such as GR effects including Shapiro delay (Backer & Hellings, 1986), gravitational redshift (Meiron et al., 2017; Torres-Orjuela et al., 2019; Chen, 2021; Yan et al., 2023), gravitational lensing (e.g. Wang et al., 1996; Takahashi & Nakamura, 2003; Ezquiaga et al., 2021; Lo et al., 2024), or contributions from tidal effects present in the triple system (see e.g. Yunes et al., 2011; Inayoshi et al., 2017; Robson et al., 2018; Chamberlain et al., 2019; Randall & Xianyu, 2019; Wong et al., 2019; Tamanini et al., 2020; D’Orazio & Loeb, 2020; Toubiana et al., 2021; Strokov et al., 2022; Xuan et al., 2023; Laeuger et al., 2023; Tiwari et al., 2024), will not be included in the present study.
@@
Figure 2: Illustration of our setup. Top: the turquoise curve represents the real trajectory of the COM of an eccentric binary (BH1, BH2), whose COM is on an eccentric orbit around a perturber (BH3). In orange, we show the trajectory of the same binary, if it had evolved in isolation. The reference binary has a constant speed $v_{m}$ equal to the orbital speed of the COM of the perturbed binary at merger, which happens at true anomaly $f_{m}$. A Rømer delay $\Delta t$ arises due to the difference in light travel time between the two scenarios towards an observer, who is depicted in the bottom right. Bottom: schematic representation of how the Rømer delay is imprinted in the gravitational waveform. The perturbed waveform receives a time-dependent Doppler shift with respect to its isolated counterpart, which we can turn into a phase shift by equating it with the orbital period of the binary (BH1, BH2) at the same point in time: $d\phi\approx 2\pi\Delta t/T$.
@@ 2.1 Rømer Time Delay
Using Kepler’s Laws, one can use geometrical arguments and standard mechanics to derive the Rømer time delay and corresponding phase shift, as we illustrate in the following. The solution to the case where the outer orbit is circular was presented in Samsing et al. (2024), and we therefore go right into presenting the solution for the RD in the more general case of an eccentric outer orbit.
@@
where $f_{m}$ is the angle at which the binary merges and $\theta(t)$ is the angle between the position vector of the turquoise binary COM at time $t$ and its position vector at merger (see Fig. 2). Another important quantity to note is the distance between the BBH and the third object, $r(f)$, which is a function of $f$ (or equivalently, $\theta$ or time) as given by,
$r(f)=\frac{a_{\mathrm{out}}(1-e_{\mathrm{out}}^{2})}{1+e_{\mathrm{out}}\cos f}.$ (3)
With this set of equations we are now in a position to calculate the distance $l$ between the reference binary undergoing a straight line motion with velocity $v_{m}$, and the perturbed binary moving on the eccentric outer orbit. Using standard geometry the distance $l$ is at time $t$ given by,
@@
where the quantities $v_{m,x}$ and $v_{m,y}$ are the $x,y$ velocity components at merger. The reduced mass $\mu$ is given by $\mu=M_{3}/(M_{1}+M_{2}+M_{3})$. These relate to the angular and radial velocities of the binary COM at phase $f$,
$v_{\text{ang}}(f)=\frac{\sqrt{GMa_{\text{out}}(1-e_{\text{out}}^{2}))}}{r(f)},$ (5)
@@
$v_{\text{rad}}(f)=e_{\text{out}}\sin(f)\sqrt{\frac{GMa_{\text{out}}}{(1-e_{\text{out}}^{2})}},$ (6)
@@
Because the outer orbit is eccentric, we cannot obtain an analytical relation between $f$ and $t$. Therefore, we calculate $l(t)$ numerically. From this calculation of the distance $l$ between the two trajectories, the RD is now simply given by,
$\Delta t(t)=l(t)/c,$ (10)
where $c$ is the speed of light.
@@ 2.2 Gravitational Wave Phase shift
As shown in Samsing et al. (2024), the maximum GW phase shift $d\phi$ evaluated in the observer frame can be approximated by the maximum RD, $\Delta{t}$, divided by the inner orbital time of the BBH, $T_{12}$, times $2\pi$ to get it in radians:
$d\phi\approx 2\pi\frac{\Delta t(t)}{T_{12}(t)}=\frac{2\pi}{c}\frac{l(t)}{T_{12}(t)}$ (11)
@@
To derive $d\phi$ for outer eccentric orbits we have the RD $\Delta{t}$ from Eq. 4 that can be evaluated at time $t$. For the corresponding orbital time $T_{12}$, given by,
$T_{12}(e)=\frac{2\pi}{\sqrt{G(M_{1}+M_{2})}}\big{(}a(e)\big{)}^{3/2},$ (12)
we need to take into account the evolution of $a$ and $e$ due to GW angular momentum and energy losses of the binary as it spirals in with time $t$. A relation that accurately describes the evolution $a(e)$ is given in Peters (1964) (which, hereafter, we will refer to as Peters64) and given by
@@
and $a_{0}$ and $e_{0}$ are, respectively, the the semi-major axis and eccentricity of the binary at some reference point in the evolution.
Note that in order to calculate the phase shift we need $\Delta t(t)$ and $T_{12}(e)$ both in terms of either $t$ or $e$. While a closed form for $e(t)$ does not exist, an expression for $t(e)$ is given in Peters64. However, this form is somewhat inaccurate for estimating the correct inspiral time when the binary is highly eccentric. The general problem relates to the fact that the orbital elements $a$ and $e$ are defined using Newtonian prescriptions, which break down especially for eccentric BBH inspirals where these quantities not only are drastically changing as a function of time, but also ill-defined. Instead we here make use of the fitting function presented in (Zwick et al., 2020, 2021), which provides a more accurate estimate for the merger time than the Peters64 solution, as a function of the initial semi-major axis $a_{0}$ and eccentricity $e_{0}$. The fit takes the form
$\begin{split}t_{c}&=\frac{5c^{5}(1+q)^{2}}{256G^{3}m^{3}q}\frac{\big{(}a(e)\big{)}^{4}}{f(e)}8^{1-\sqrt{1-e}}\exp\left(\frac{2.8r_{S}}{a(e)(1-e)}\right)\\ &\quad\times\Bigg{\{}1+\bigg{[}-1+\exp\left(\frac{2.2r_{S}}{a(e)(1-e)}\right)\\ &\qquad\quad+\left(\frac{3.8r_{S}}{a(e)(1-e)}\right)^{3/2}\bigg{]}(1-e)^{2}\\ &\qquad\quad-\left(\frac{3.8r_{S}}{a(e)(1-e)}\right)^{3/2}(1-e)\Bigg{\}},\end{split}$ (15)
where $m=M_{1}+M_{2}$, $q=M_{2}/M_{1}\geq 1$, $r_{S}=\frac{2G(M_{1}+M_{2})}{c^{2}}$ and
@@
The relation $a(e)$ from Peters64 given by Eq. 13 has been shown to still provide an accurate description, and we therefore continue using this throughout this work.

### v2

@@ Abstract
We study the gravitational wave (GW) phase shift arising from center-of-mass accelerations of binary black hole mergers formed dynamically in three-body systems, where both the inner orbit of the merging binary and the outer orbit are eccentric. We provide a semi-analytical model and several analytical approximations that allow for fast evaluation of both the temporal evolution and the maximum value of the phase shift. The highest phase shifts occur when the binary merges close to the pericentre of the outer orbit, and can in this case be orders-of-magnitude larger compared to the circular limit. At high outer orbit eccentricities, the orbital curvature leaves distinct imprints onto the phase shift if the binary passes the outer pericentre during its inspiral. By comparing with phase shifts measured in numerical chaotic 3-body scatterings, we show that our model accurately describes the observed phase of dynamically assembled binary systems in realistic astrophysical scenarios, providing a way to directly determine their formation channel via single GW observations.
@@ 1 Introduction
Figure 1: Formation of a GW phase-shifted source. Illustrative example of a scattering between a binary black hole (BH1,BH3) and an incoming single black hole (BH2) that results in a highly eccentric BBH merger (BH1,BH2), while all three objects are still bound to each other. The eccentric BBH merger therefore inspirals and merge on an eccentric orbit around the remaining BH3, which gives rise to a Doppler-shifted GW signal due to binary COM acceleration in the COM frame. This GW-shift can be detected, and will leave imprints on how the specific BBH assembly and merger takes place, thereby revealing the BBH origin. This is further illustrated in Fig. 2. In this example, we use $m_{1}=15M_{\odot}$, $m_{2}=5M_{\odot}$, and $m_{3}=15M_{\odot}$.
One very interesting measure of the formation environment is how the presence of a third body induces a phase shift onto the gravitational waveform in the observer frame. In such a scenario, the BBH evolves while its COM is being accelerated. The acceleration causes a time-dependent Doppler shift in the gravitational waveform, which in principle can be measured and used to probe the exact dynamical environment within which the binary lives as it merges. Therefore, looking for GW phase shifts, or more generally environmental effects, could be the key to help us tell apart the large suit of sub-channels, provided they are present in the observable GW bands (e.g. Tiwari et al., 2024b, a).
@@
This generalisation is vital as it accurately represents the case for dynamically assembled BBH mergers (see e.g., Portegies Zwart & McMillan, 2000; Samsing et al., 2014b; Rodriguez et al., 2016a, b; Askar et al., 2017; Samsing & D’Orazio, 2018; Liu & Lai, 2018; Hoang et al., 2018; Trani et al., 2022). An example is depicted in Fig. 1, that shows the formation of a BBH inspiral and merger formed through a chaotic scattering process between three black holes. These types of interactions are known to frequently take place in GCs and other dense environments, and have especially been suggested to be among the most reliable ways for forming BBH mergers with residual eccentricity across the observable GW bands (Samsing, 2018). Moreover, including and considering outer orbit eccentricity is not expected to just add a small perturbation to the problem, as the maximum GW phase shift scales as $\propto R^{-2}$, where $R$ is the distance between the BBH and the third BH (Meiron et al., 2017; Samsing et al., 2024). This implies that if the BBH merges close to the pericentre of its outer orbit, the GW phase shift may increase dramatically compared to a circular outer orbit with the same period. This situation is in fact seen in Fig. 1, where ‘merger’ takes place near pericentre of the outer orbit.
The aim of this paper is to investigate the GW Doppler shift and corresponding GW phase shift of an eccentric BBH on an eccentric orbit around a third BH, by quantifying effects that are unique to the outer orbit being eccentric. Effects for different outer orbit families have recently been explored using an expansion in acceleration (Tiwari et al., 2024a); In our present paper we expand on this, by further developing a semi-analytical method that solves for the full orbit from which we discover several new non-linear effects in the resultant GW phase shift.
@@ 2 Theory
As described in Meiron et al. (2017); Samsing et al. (2024), the GW Doppler shift arising from the acceleration on an outer orbit can equivalently be viewed as the GW phase shift from Rømer delay (RD) between the binary trajectory and some non-accelerating reference trajectory. In Fig. 2 we show our setup in this framework. We depict in turquoise the trajectory of an initially highly (inner-orbit) eccentric BBH (BH1, BH2) that inspirals and merges on an eccentric (outer) orbit around BH3 (see also Fig. 1). Each of these BHs have a mass $m$ denoted $m_{1}$, $m_{2}$ and $m_{3}$, respectively. We compare this to a fictional reference binary, shown in orange. This binary has exactly the same parameters (i.e. initial eccentricity, semi-major axis, time to coalescence, binary masses), but moves on a straight line with a constant velocity $v_{m}$ that is equal to the tangential velocity of the perturbed binary at merger. Tracing these two paths back in time from merger, their gradual deviation is what gives rise to the time-dependent Rømer delay. If the distance between the two trajectories at some time $t$ is given by $l(t)$ (red dashed line in Fig. 2), then their relative maximum Rømer delay, $\Delta t(t)$, is
$\Delta t(t)=l(t)/c,$ (1)
where $c$ is here the speed of light. As shown in the bottom sketch of Fig. 2, the difference in arrival time between the GWs from the two trajectories induces a phase shift onto the gravitational waveform.
It is important to realise that the $\Delta t$ shown in Fig. 2 is the maximum time delay between the two trajectories. In reality, an observer only sees the projection of $\Delta t$ in their own line-of-sight (e.g. Meiron et al., 2017; Samsing et al., 2024). However, in this work, we solely focus on the maximum possible phase shift that we denote simply by $d\phi$; factoring in the trivial observer dependent factor will be investigated in later work. Other sources of dephasing in gravitational waveforms in dynamically assembled systems includes General Relativistic (GR) effects, such as Shapiro delay (Backer & Hellings, 1986), gravitational redshift (Meiron et al., 2017; Chen, 2021), and gravitational lensing (e.g. Wang et al., 1996; Takahashi & Nakamura, 2003; Ezquiaga et al., 2021; Lo et al., 2024), as well as special relativistic effects (Torres-Orjuela et al., 2019; Yan et al., 2023), aberration (Torres-Orjuela et al., 2020), and classical tidal effects (Samsing et al., 2024). Relevant discussions of similar setups can be found in (e.g. Yunes et al., 2011; Inayoshi et al., 2017; Robson et al., 2018; Chamberlain et al., 2019; Randall & Xianyu, 2019; Wong et al., 2019; Tamanini et al., 2020; D’Orazio & Loeb, 2020; Toubiana et al., 2021; Strokov et al., 2022; Xuan et al., 2023; Laeuger et al., 2023; Tiwari et al., 2024b).
@@
Figure 2: Illustration of our setup. Top: The turquoise curve represents the real trajectory of the COM of an eccentric binary (BH1, BH2), whose COM is on an eccentric orbit around a perturber (BH3). In orange, we show the trajectory of the same binary, if it had evolved in isolation. The reference binary has a constant speed $v_{m}$ equal to the orbital speed of the COM of the perturbed binary at merger, which happens at true anomaly $f_{m}$. A Rømer delay $\Delta t$ arises due to the difference in light travel time between the two scenarios towards an observer, who is depicted in the bottom right. Bottom: schematic representation of how the Rømer delay is imprinted in the gravitational waveform. The perturbed waveform receives a time-dependent Doppler shift with respect to its isolated counterpart, which we can turn into a phase shift by equating it with the orbital period of the binary (BH1, BH2) at the same point in time: $d\phi\approx 2\pi\Delta t/T_{12}$.
@@ 2.1 Rømer Time Delay
Using Kepler’s Laws, one can use geometrical arguments and standard mechanics to derive the Rømer time delay and corresponding phase shift, as we illustrate in the following. The analytical solution to the case where the inner orbit is eccentric and the outer orbit is circular was presented in Samsing et al. (2024), and we therefore go right into presenting the solution for the RD in the more general case of an eccentric outer orbit.
@@
where $f_{m}$ is the angle at which the binary merges and $\theta(t)$ is the angle between the position vector of the turquoise binary COM at time $t$ and its position vector at merger (see Fig. 2). Another important quantity to note is the distance between the BBH and the third object, $r(f)$, which is a function of $f$ as given by,
$r(f)=\frac{a_{\mathrm{out}}(1-e_{\mathrm{out}}^{2})}{1+e_{\mathrm{out}}\cos f},$ (3)
where $a_{\mathrm{out}}$ and $e_{\mathrm{out}}$ denote the semi-major axis and eccentricity of the BBH outer orbit, respectively. With this set of equations we are now in a position to calculate the distance $l$ between the reference binary undergoing a straight line motion with velocity $v_{m}$, and the perturbed or true binary moving on the eccentric outer orbit. Using standard geometry the distance $l$ is at time $t$ given by,
@@
where the quantities $v_{m,x}$ and $v_{m,y}$ are the $x,y$ velocity components at merger, $f$ depends on $t$ through Eq. 2, and $\mu=m_{3}/m_{123}$, where $m_{123}=m_{1}+m_{2}+m_{3}$. In the rest of the paper, we will refer to the quantity $r(f_{m})$, i.e. the distance to the perturber at merger, simply as $r_{m}$. These quantities relate to the angular and radial velocities of the binary COM at phase $f$,
$v_{\text{ang}}(f)=\frac{\sqrt{Gm_{123}a_{\text{out}}(1-e_{\text{out}}^{2}))}}{r(f)},$ (5)
@@
$v_{\text{rad}}(f)=e_{\text{out}}\sin(f)\sqrt{\frac{Gm_{123}a_{\text{out}}}{(1-e_{\text{out}}^{2})}},$ (6)
@@
Because the outer orbit is eccentric, we cannot obtain an analytical relation between $f$ and $t$. Therefore, we calculate $l(t)$ numerically using the equations above. From this calculation of the distance $l$ between the two trajectories, the RD can now be calculated using Eq. 1.
@@ 2.2 Gravitational Wave Phase shift
As shown in Samsing et al. (2024), the maximum GW phase shift at a given time $t$, $d\phi(t)$, evaluated in the observer frame can be approximated by the maximum RD, $\Delta{t}$, divided by the inner orbital time of the BBH, $T_{12}$, times $2\pi$ to get it in radians,
$d\phi(t)\approx 2\pi\frac{\Delta t(t)}{T_{12}(t)}=\frac{2\pi}{c}\frac{l(t)}{T_{12}(t)}.$ (10)
@@
To derive $d\phi$ for outer eccentric orbits we have the RD $\Delta{t}$ from Eq. 4 that can be evaluated at time $t$. The corresponding orbital time $T_{12}$ can also be defined at time $t$, which follows from Kepler’s Law,
$T_{12}=\frac{2\pi}{\sqrt{G(m_{1}+m_{2})}}a^{3/2},$ (11)
where $a$, and associated $e$, refer to the semi-major axis and the eccentricity of the merging BBH inner orbit, respectively. For solving the evolution, from assembly to merger of the BBH due to GW angular momentum and energy losses as it spirals in, we use the relation for $a(e)$ presented in Peters (1964) (which, hereafter, we will refer to as Peters64),
@@
and $a_{0}$ and $e_{0}$ are the semi-major axis and eccentricity of the binary at some reference point in their evolution.
Note that in order to calculate the phase shift we need $\Delta t(t)$ and $T_{12}$ in terms of either $t$ or $e$. While a closed form for $e(t)$ does not exist, an expression for $t(e)$ is given in Peters64. However, this form is somewhat inaccurate for estimating the correct inspiral time when the binary is highly eccentric. The general problem relates to the fact that the orbital elements, $a$ and $e$, are defined using Newtonian prescriptions, which break down especially for eccentric BBH inspirals where these quantities not only are drastically changing as a function of time, but also ill-defined. Instead we here make use of the fitting function presented in (Zwick et al., 2020, 2021), which provides a more accurate estimate for the merger time than the Peters64 solution. The fit takes the form
$\begin{split}t_{c}&=\frac{5c^{5}(1+q)^{2}}{256G^{3}m_{12}^{3}q}\frac{a(e)^{4}}{f(e)}8^{1-\sqrt{1-e}}\exp\left(\frac{2.8r_{S}}{a(e)(1-e)}\right)\\ &\quad\times\Bigg{\{}1+\bigg{[}-1+\exp\left(\frac{2.2r_{S}}{a(e)(1-e)}\right)\\ &\qquad\quad+\left(\frac{3.8r_{S}}{a(e)(1-e)}\right)^{3/2}\bigg{]}(1-e)^{2}\\ &\qquad\quad-\left(\frac{3.8r_{S}}{a(e)(1-e)}\right)^{3/2}(1-e)\Bigg{\}},\end{split}$ (14)
where $m_{12}=m_{1}+m_{2}$, $q=m_{2}/m_{1}\geq 1$, $r_{S}={2G(m_{1}+m_{2})}/{c^{2}}$ and
@@
The relation $a(e)$ from Peters64 given by Eq. 12, has been shown to still provide an accurate description, and we therefore continue using this throughout this work.